forty divided by a number is 8
Answer by Theo(13342)

let x be the number
40/x = 8
multiply both sides of the equation by x to get:
40 = 8x
divide both sides of the equation by 8 to get:
40/8 = x
solve for x to get:
x = 5
40/5 = 8